# On computing and the complexity of computing higher-order $U$-statistics, exactly

### v1

@@ 2.1 UU/$V$-Statistics and Tensors
In statistical applications, $\mathbb{X}$} and the tuple $\bm{X}$, respectively, correspond to the sample space and the observed data sample of size $n$. Unlike $U$-statistics, computing $V$-statistics amounts to summing over all $n]^{m}$ without being restricted to permutations. As will be clear later in this paper, $V$-statistics can be represented as summations over higher-order tensors. This unique property renders computing $V$-statistics, but not $U$-statistics, amenable for tensor operations such as the Einstein summation (einstein1916grundlage). In particular, if the kernel function $h$ satisfies certain decomposable structure (see Definition 9 later), the computational complexity of $V$-statistics can be notably reduced. Our algorithmic framework is thus to first turn a $U$-statistic into a linear combination of $V$-statistics and then compute each individual $V$-statistic via the Einstein summation.
2.2 The Einsum Notation and Einsum Operation
As mentioned, our algorithm to be revealed will rely on the Einstein summation operation (abbreviated as the Einsum operation) over tensors. In this section, we give a brief but (hopefully) pedagogical overview for readers unfamiliar with these related concepts. We also steer readers to klaus2023compiling or blacher2024einsum for more details.
The Einsum operation to be defined serves as a fundamental computational primitive in our proposed algorithm for efficiently and exactly computing UU/$V$-statistics. The Einsum operation generalizes the classical Einstein summation convention introduced in (einstein1916grundlage). The modern Einsum operation, now implemented in popular Python packages such as numpy and pytorch, can be used for efficient tensor computation. To explain the Einsum operation, we first introduce a concept called the Einsum notation.
Definition 4 (The Einsum Notation).
@@ 2.2 The Einsum Notation and Einsum Operation
${\mathcal{A}}\coloneqq(A_{k})_{k=1}^{K}$ is referred to as the input tuple list of $\mathcal{N}$} and $B$ is referred as the output tuple. The index set of $\mathcal{N}$} is the set of all indices appearing in the input tuple list, and is denoted by $\mathsf{Indices}({\mathcal{N}})$. Without loss of generality, we always take $m$ to be the smallest $m$ such that $m]=\mathsf{Indices}({\mathcal{N}})$. If $B=\emptyset$, we then simply write $\mathcal{A}$} as ${\mathcal{N}}=({\mathcal{A}};\emptyset)$ and call $\mathcal{A}$} as the Einsum notation.
It is noteworthy that in the above definition, the index set $[m]$ is simply chosen for notational convenience and in principle the indices do not have to be integers or even numbers. We now define the Einsum operation based on the Einsum notation.
Definition 5 (The Einsum Operation).
Let ${\mathcal{N}}=({\mathcal{A}};B)$ be an Einsum notation w.r.t. the index set $I=[m]$. Suppose that a set of tensors ${\mathcal{T}}\coloneqq(T_{k})_{k=1}^{K}$, each of size $n$ and of order $(m_{k})_{k=1}^{K}$, can be represented by $\mathcal{N}$}, i.e. $m_{k}=|A_{k}|$ for any $k\in[K]$. An Einsum operation, denoted as $\textsf{Einsum}({\mathcal{T}},{\mathcal{N}})$, takes $\mathcal{T}$} and $\mathcal{N}$} as inputs and outputs another tensor $T^{\rm out}:[n]^{|B|}\rightarrow{\mathbb{R}}$ as follows: for any $\alpha\in[n]^{|B|}$,
@@
We now use several concrete examples to illustrate the above definitions.
Example 1.
We take matrix multiplication as an example to illustrate the Einsum operation just defined. Let $T_{1},T_{2}$ be two input matrices, which are also 2nd-order tensors. Therefore ${\mathcal{T}}=(T_{1},T_{2})$. Suppose we are interested in multiplying $T_{1}$ and $T_{2}$, denoted as $T^{\rm out}$: for any $i,j\in[n]$,
$T^{\rm out}(i,j)=\sum_{k=1}^{n}T_{1}(i,k)T_{2}(k,j).$
The matrix multiplication corresponds to the Einsum operation with Einsum notation ${\mathcal{N}}=({\mathcal{A}};B)$, with input tuple list ${\mathcal{A}}=((1,2),(2,3))$ and output tuple $B=(1,3)$ and $|B|=2$. The index set is simply $I=[3]$. This means that the operation sums over the dimensions of $T_{1}$ and $T_{2}$ corresponding to index $2$ and reserves indices 11 and $3$. $T^{\rm out}(\alpha)$ for any $\alpha=(i,j)\in[n]^{2}$ can be written in the Einsum operation form $\textsf{Einsum}({\mathcal{T}},{\mathcal{N}})$ because by definition:
$\textsf{Einsum}({\mathcal{T}},{\mathcal{N}})$ $\equiv\sum_{\alpha^{\prime}\in[n]^{3}:\alpha^{\prime}[1,3]=\alpha}\prod_{k=1}^{2}T_{k}(\alpha^{\prime}[A_{k}])$
$=\sum_{\alpha^{\prime}\in[n]^{3}:\alpha^{\prime}[1,3]=\alpha}T_{1}(\alpha^{\prime}[A_{1}])T_{2}(\alpha^{\prime}[A_{2}])$
$=\sum_{\alpha^{\prime}\in[n]^{3}:\alpha^{\prime}[1,3]=\alpha}T_{1}(\alpha^{\prime}[1,2])T_{2}(\alpha^{\prime}[2,3])$
$=\sum_{\alpha^{\prime}\in[n]^{3}:\alpha^{\prime}[1,3]=\alpha}T_{1}(\alpha^{\prime}[1],\alpha^{\prime}[2])T_{2}(\alpha^{\prime}[2],\alpha^{\prime}[3])$
$=\sum_{\alpha^{\prime}[2]=1}^{n}T_{1}(\alpha[1],\alpha^{\prime}[2])T_{2}(\alpha^{\prime}[2],\alpha[3])$
$=\sum_{\alpha^{\prime}[2]=1}^{n}T_{1}(i,\alpha^{\prime}[2])T_{2}(\alpha^{\prime}[2],j),$
so the claim holds by identifying $\alpha^{\prime}[2]=k$.
Example 2.
$T^{\rm out}(i,j,k)$ is in fact a partial trace, a common operation taken in quantum mechanics. The corresponding Einsum notation is ${\mathcal{N}}=({\mathcal{A}};B)$ with ${\mathcal{A}}=(A_{1}=(1,2,3),A_{2}=(2,4,5),A_{3}=(3,6))$ with $B=(1,5,6)$ and the index set $I=[6]$. $T^{\rm out}(\alpha)$ for any $\alpha=(i,j,k)\in[n]^{3}$ can be written in the Einsum operation form $\textsf{Einsum}({\mathcal{T}},{\mathcal{N}})$ because by definition:
$\textsf{Einsum}({\mathcal{T}},{\mathcal{N}})$ $\equiv\sum_{\alpha^{\prime}\in[n]^{4}:\alpha^{\prime}[1,5,6]=\alpha}T_{1}(\alpha^{\prime}[A_{1}])T_{2}(\alpha^{\prime}[A_{2}])T_{3}(\alpha^{\prime}[A_{3}])$
$=\sum_{\alpha^{\prime}\in[n]^{4}:\alpha^{\prime}[1,5,6]=\alpha}T_{1}(\alpha^{\prime}[1,2,3])T_{2}(\alpha^{\prime}[2,4,5])T_{3}(\alpha^{\prime}[3,6])$
$=\sum_{\alpha^{\prime}\in[n]^{4}:\alpha^{\prime}[1,5,6]=\alpha}T_{1}(\alpha^{\prime}[1,2,3])T_{2}(\alpha^{\prime}[2,4,5])T_{3}(\alpha^{\prime}[3,6])$
$=\sum_{\alpha^{\prime}[2]=1}^{n}\sum_{\alpha^{\prime}[3]=1}^{n}\sum_{\alpha^{\prime}[4]=1}^{n}T_{1}(\alpha[1],\alpha^{\prime}[2],\alpha^{\prime}[3])T_{2}(\alpha^{\prime}[2],\alpha^{\prime}[4],\alpha[5])T_{3}(\alpha^{\prime}[3],\alpha[6]),$
so the claim holds by identifying $\alpha^{\prime}[2]=a$, $\alpha^{\prime}[3]=b$, and $\alpha^{\prime}[4]=c$.
@@
The corresponding Einsum notation is ${\mathcal{A}}=(A_{1}=(1,2),A_{2}=(2,3),A_{3}=(3,4))$ with $B=\emptyset$ and the index set $I=[4]$. $T^{\rm out}$ can be written in the Einsum operation form $\textsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ because by definition:
$\textsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ $\equiv\sum_{\alpha\in[n]^{4}}\prod_{k=1}^{3}T_{k}(\alpha[A_{k}])=\sum_{\alpha\in[n]^{4}}T_{1}(\alpha[A_{1}])T_{2}(\alpha[A_{2}])T_{3}(\alpha[A_{3}])$
@@
The Einsum operation has been implemented in highly optimized Python libraries such as numpy.einsum and pytorch.einsum. These efficient algorithmic subroutines, enabling efficient Einsum operation in practice, serve as the building blocks for our algorithm. Notably, developed for training large overparameterized neural networks in tensor forms, pytorch.einsum is also designed to leverage GPU computing to further enhance the scalability and speed of the Einsum operation, as we will demonstrate using an example from statistics in Section 4.1.
@@
The Einsum operation is frequently employed in scientific computing problems. For example, the Einsum operation serves as a basic algorithmic subroutine for computation over tensor (hyper)networks, a useful mathematical language adopted by the quantum computing and quantum information community (markov2008simulating; stoudenmire2018learning). Although we do not directly adopt the tensor network formulation, our strategy is quite similar in that we also establish a correspondence between the Einsum operation and operations on graphs – see Section 3.2 for more details.
@@ 2.3 Useful Concepts from Graph Theory
In this section, we recall three key concepts from graph theory that are central to our complexity-theoretic characterization: vertex elimination and treewidth, which captures the dominant computational complexity; and quotient graph, which is related to the decomposition of a $U$-statistic into a linear combination of $V$-statistics. Here we adopt the terminologies from bodlaender2010treewidth.
@@
3 A New Algorithm of Exactly Computing $U$-Statistics
We are ready to present our new algorithm for exactly computing $U$-statistics, along with an analysis of the time complexity. Our algorithm has been implemented in an open source Python package u-stats.
We begin by introducing the multiplicative-decomposable (MD) structure of the UU/$V$-statistic kernel in Section 3.1. We then describe our algorithm and the complexity analysis for $V$-statistics in Section 3.2, followed by those for $U$-statistics in Section 3.3.
@@ 3.1 The Multiplicative-Decomposable Kernels
Definition 9 (Multiplicative Decomposition of a Function).
Let $\mathbb{X}$} be a non-empty set, $m$ a positive integer, and $h:{\mathbb{X}}^{m}\to{\mathbb{R}}$. Suppose ${\mathcal{A}}\coloneqq(A_{k})_{k=1}^{K}\in[m]^{**}$ satisfies $\bigcup_{k=1}^{K}\mathsf{set}(A_{k})=[m]$, and let ${\mathcal{H}}\coloneqq(h_{k})_{k=1}^{K}$ be a tuple of functions where each $h_{k}$ is defined on ${\mathbb{X}}^{|A_{k}|}$. ${\mathcal{H}}\times{\mathcal{A}}$, comprised of ordered pairs of the form $(h_{k},A_{k})$ for $k\in[K]$, is called a multiplicative decomposition of the function $h$, if for any $\bm{x}\in{\mathbb{X}}^{m}$,
@@
In the sequel, we refer to the function tuple $\mathcal{H}$} as the kernel components and the index tuple $\mathcal{A}$} as the decomposition signature. For convenience, we say that a kernel $h$ permits a (nontrivial) MD structure ${\mathcal{H}}\times{\mathcal{A}}$ when $K>1$.
$U$-statistic kernels with MD structures are pervasive in statistical applications. We will provide three examples in Section 4.
To utilize Einsum operation to compute UU/$V$-statistics, we need to store pieces of the underlying statistics in tensor format. We term this step as tensorization. Let $h$ be the kernel of some UU/$V$-statistic permitting a MD structure ${\mathcal{H}}\times{\mathcal{A}}$, where ${\mathcal{H}}\coloneqq(h_{k})_{k=1}^{K}$ and ${\mathcal{A}}\coloneqq(A_{k})^{K}_{k=1}$. Given a sample $\bm{X}\in{\mathbb{X}}^{n}$, the tensorization step constructs a tuple of tensors ${\mathcal{T}}\coloneqq(T_{k})_{k=1}^{K}$ as follows:
@@
Then following Definition 2 and Definition 9, we have
@@
4:A tuple of tensors ${\mathcal{T}}\coloneqq(T_{k}:[n]^{m_{k}}\to{\mathbb{R}})_{k=1}^{K}$, where $T_{k}$ is the tensorized version of $h_{k}$
@@
$\mathbb{V}(h,\bm{X})=\mathbb{V}({\mathcal{T}},{\mathcal{A}})=\textsf{Einsum}{}({\mathcal{T}},{\mathcal{A}}),\quad\mathbb{U}(h,\bm{X})=\mathbb{U}({\mathcal{T}},{\mathcal{A}}).$
3.2 The Algorithm and Time Complexity of Computing $V$-Statistics
@@
$\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}})=\sum_{\alpha\in[n]^{m}}\prod_{k=1}^{K}T_{k}(\alpha[A_{k}]).$
@@
$\mathbb{V}(h,\bm{X})=\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}}),$
@@
7:  $v\leftarrow\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ [^1] In our implementation of u-stats, we directly call numpy.einsum or pytorch.einsum, with the summation ordering optimized by opt-einsum.
@@
The time complexity of computing a $V$-statistic can obviously be characterized by that of the Einsum operation. The essential idea is to establish a correspondence between the Einsum operation and the vertex elimination operation (Definition 6) over a graph associated with the Einsum notation. The time complexity can in turn be characterized by the treewidth (Definition 7) of that graph. Similar idea has also been exploited in markov2008simulating, in which the time complexity of as the so-called tensor contraction operation (a special case of the Einsum operation) is also characterized via the treewidth of the underlying tensor network.
@@
Definition 10 (The Decomposition Graph Associated with an Einsum Notation).
@@
  • $V_{{\mathcal{A}}}=\mathsf{Indices}({\mathcal{A}})$,
@@
Example 4 (Correspondence between the Einsum notation and the decomposition graph).
@@
Figure 1: Examples of decomposition graphs associated with different decomposition signatures.
We are ready to present the following “optimistic” estimate of the time complexity of the Einsum operation. The proof of the proposition below is deferred to Appendix A.
@@
Let $\mathcal{A}$} be an Einsum notation with no output, and $G_{\mathcal{A}}$ be the corresponding decomposition graph of $\mathcal{A}$}. Then there exists an algorithm with a particular Einsum ordering such that for any set of tensors $\mathcal{T}$} that can be represented by the Einsum notation $\mathcal{A}$}, the time complexity of $\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ is $O(|{\mathcal{A}}|n^{\mathsf{tw}(G_{\mathcal{A}})+1})$.
Remark 4 (The Einsum ordering in practice).
We say that the time complexity given in the above proposition is only an optimistic estimate, because although it can be achieved with the optimal Einsum ordering corresponding to the treewidth (Definition 7), this particular ordering is usually difficult to find in practice. arnborg1987complexity have shown that finding the ordering that gives rise to the treewidth is an NP-complete problem. While it is ideal to evaluate $\textsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ in Algorithm 2 using this optimal ordering, we decide to directly plug in the Python library opt-einsum (daniel2018opt) to find the ordering in practice, because such a preexisting library has been optimized and stabilized over the years. opt-einsum offers both heuristic strategies such as greedy algorithms or exhaustive search such as dynamic programming, but the latter can be time-consuming when dealing with large-scale problems. As mentioned in the Introduction, dynamic programming is also used in zhang2025adaptive for a special class of $U$-statistics. Since the ordering only depends on the MD structure of the kernel $h$, for commonly encountered kernels, one could in principle find the optimal ordering via exhaustive search without accessing the data sample and re-use this ordering whenever the same type of kernel appears. Finally, we briefly remark that opt-einsum uses a different scheme to describe the summation ordering from the one used in this paper. We will not describe such a difference in any detail as doing so will necessarily introduce extra notations and definitions that are only tangential to the main theme of the paper.
@@
3.3 The Algorithm and Time Complexity of Computing $U$-Statistics

### v2

@@ 2.1 UU/$V$-Statistic and Tensor
In statistical applications, $\mathbb{X}$ and the tuple $\bm{X}$, respectively, correspond to the sample space and the observed data sample of size $n$. Unlike $U$-statistics, computing $V$-statistics amounts to summing over all $n]^{m}$ without being restricted to permutations. As will be clear later in this paper, $V$-statistics can be represented as summations over higher-order tensors. This unique property renders computing $V$-statistics, but not $U$-statistics, amenable for tensor operations such as the Einstein summation. In particular, if the kernel function $h$ satisfies certain decomposable structure (see Definition 9 later), the computational complexity of $V$-statistics can be notably reduced. Our algorithmic framework is thus to first turn a $U$-statistic into a linear combination of $V$-statistics and then compute each individual $V$-statistic via the Einstein summation.
2.2 Einsum Notation and Einsum Operation
As mentioned, our algorithm to be revealed will rely on the Einstein summation operation (abbreviated as the Einsum operation) over tensors. In this section, we give a brief but (hopefully) pedagogical overview for readers unfamiliar with these related concepts. We also steer readers to blacher2024einsum for more details.
The Einsum operation to be defined serves as a fundamental computational primitive in our proposed algorithm to efficiently and exactly compute the UU/$V$-statistics. The Einsum operation generalizes the classical Einstein summation convention. The modern Einsum operation, now implemented in popular Python packages such as numpy and pytorch, can be used for efficient tensor computation. To explain the Einsum operation, we first introduce a concept called the Einsum notation.
Definition 4 (Einsum Notation).
@@ 2.2 Einsum Notation and Einsum Operation
$\mathcal{A}\coloneqq(A_{k})_{k=1}^{K}$ is referred to as the input tuple list of $\mathcal{N}$ and $B$ is referred as the output tuple. The index set of $\mathcal{N}$ is the set of all indices appearing in the input tuple list, and is denoted by $\textsc{Indices}(\mathcal{N})$. Without loss of generality, we always take $m$ to be the smallest $m$ such that $m]=\textsc{Indices}(\mathcal{N})$. If $B=\emptyset$, we then simply write $\mathcal{A}$ as $\mathcal{N}=(\mathcal{A};\emptyset)$ and call $\mathcal{A}$ as the Einsum notation.
It should be noted that, in the above definition, the index set $[m]$ is simply chosen for notational convenience and, in principle, the indices need not be integers or numbers. We now define the Einsum operation based on the Einsum notation.
Definition 5 (Einsum Operation).
Let $\mathcal{N}=(\mathcal{A};B)$ be an Einsum notation w.r.t. the index set $I=[m]$. Suppose that a set of tensors $\mathcal{T}\coloneqq(T_{k})_{k=1}^{K}$, each of size $n$ and of order $(m_{k})_{k=1}^{K}$, can be represented by $\mathcal{N}$, i.e. $m_{k}=|A_{k}|$ for any $k\in[K]$. An Einsum operation, denoted as $\textsc{Einsum}(\mathcal{T},\mathcal{N})$, takes $\mathcal{T}$ and $\mathcal{N}$ as inputs and outputs another tensor $T^{\rm out}:[n]^{|B|}\rightarrow{\mathbb{R}}$ as follows: for any $\alpha\in[n]^{|B|}$,
@@
[…]
@@
The corresponding Einsum notation is $\mathcal{A}=(A_{1}=(1,2),A_{2}=(2,3),A_{3}=(3,4))$ with $B=\emptyset$ and the index set $I=[4]$. $T^{\rm out}$ can be written in the Einsum operation form $\textsc{Einsum}(\mathcal{T},\mathcal{A})$ because by definition:
$\textsc{Einsum}(\mathcal{T},\mathcal{A})$ $\equiv\sum_{\alpha\in[n]^{4}}\prod_{k=1}^{3}T_{k}(\alpha[A_{k}])=\sum_{\alpha\in[n]^{4}}T_{1}(\alpha[A_{1}])T_{2}(\alpha[A_{2}])T_{3}(\alpha[A_{3}])$
@@
The Einsum operation has been implemented in highly optimized Python libraries such as numpy.einsum and pytorch.einsum. These efficient algorithmic subroutines, enabling efficient Einsum operation in practice, serve as the building blocks of our algorithm. Notably, developed for training large overparameterized neural networks in tensor forms, pytorch.einsum is also designed to leverage GPU computing to further enhance the scalability and speed of the Einsum operation, as we will demonstrate using an example from statistics in Section 4.1.
@@
The Einsum operation is frequently employed in scientific computing problems. For example, the Einsum operation serves as a basic algorithmic subroutine for computation over tensor (hyper)networks, a useful mathematical language adopted by the quantum computing and quantum information community (markov2008simulating). Although we do not directly adopt the tensor network formulation, our strategy is quite similar in that we also establish a correspondence between the Einsum operation and operations on graphs – see Section 3.2 for more details.
@@ 2.3 Useful Concepts from Graph Theory
In this section, we recall three key concepts from graph theory that are central to our complexity-theoretic characterization. We follow the terminology of bodlaender2010treewidth for vertex elimination and treewidth, which captures the dominant computational complexity of our problem. We also recall the notion of quotient graph, adopted from sanders2012high, which underlies the decomposition of a $U$-statistic into a linear combination of $V$-statistics.
@@
3 A New Algorithm of Computing $U$-Statistics
We are ready to present our new algorithm for exactly computing $U$-statistics, along with an analysis of the time complexity. Our algorithm has been implemented in an open source Python package u-stats (and its R interface). We begin by introducing the multiplicative-decomposable (MD) structure of the UU/$V$-statistic kernel in Section 3.1, followed by the new algorithm and complexity analysis for $V$- and $U$-statistics in Section 3.2 and Section 3.3.
@@ 3.1 The Multiplicative-Decomposable Kernels
Definition 9 (Multiplicative Decomposition).
Let $\mathbb{X}$ be a non-empty set, $m$ a positive integer, and $h:\mathbb{X}^{m}\to{\mathbb{R}}$. Suppose $\mathcal{A}\coloneqq(A_{k})_{k=1}^{K}\in[m]^{**}$ satisfies $\bigcup_{k=1}^{K}\mathsf{set}(A_{k})=[m]$, and let $\mathcal{H}\coloneqq(h_{k})_{k=1}^{K}$ be a tuple of functions where each $h_{k}$ is defined on $\mathbb{X}^{|A_{k}|}$. $\mathcal{H}\times\mathcal{A}$, consisting of ordered pairs of the form $(h_{k},A_{k})$ for $k\in[K]$, is called a multiplicative decomposition of the function $h$, if for any $\bm{x}\in\mathbb{X}^{m}$,
@@
[…]
To utilize the Einsum operation to compute UU/$V$-statistics, we need to store pieces of the underlying statistics in tensor format. We term this step as tensorization. Let $h$ be the kernel of some UU/$V$-statistic permitting a MD structure $\mathcal{H}\times\mathcal{A}$, where $\mathcal{H}\coloneqq(h_{k})_{k=1}^{K}$ and $\mathcal{A}\coloneqq(A_{k})^{K}_{k=1}$. Given a sample $\bm{X}\in\mathbb{X}^{n}$, the tensorization step constructs a tuple of tensors $\mathcal{T}\coloneqq(T_{k})_{k=1}^{K}$ as follows:
@@
Then following Definition 2 and Definition 9,
@@
4:A tuple of tensors $\mathcal{T}\coloneqq(T_{k}:[n]^{m_{k}}\to{\mathbb{R}})_{k=1}^{K}$, where $T_{k}$ is the tensorized $h_{k}$
@@
$\mathbb{V}(h,\bm{X})=\mathbb{V}(\mathcal{T},\mathcal{A})=\textsc{Einsum}{}(\mathcal{T},\mathcal{A}),\quad\mathbb{U}(h,\bm{X})=\mathbb{U}(\mathcal{T},\mathcal{A}).$
3.2 The Algorithm of Computing $V$-Statistics
@@
$\textsc{Einsum}(\mathcal{T},\mathcal{A})=\sum_{\alpha\in[n]^{m}}\prod_{k=1}^{K}T_{k}(\alpha[A_{k}]).$
@@
$\mathbb{V}(h,\bm{X})=\textsc{Einsum}(\mathcal{T},\mathcal{A}),$
@@
7:  $v\leftarrow\textsc{Einsum}(\mathcal{T},\mathcal{A})$ [^1] In our implementation of u-stats, we directly call numpy.einsum or pytorch.einsum, with the summation ordering optimized by opt-einsum.
@@
The time complexity of computing a $V$-statistic can obviously be characterized by that of the Einsum operation. The essential idea is to establish a correspondence between the Einsum operation and the vertex elimination operation (Definition 6) over a graph associated with the Einsum notation. The time complexity can in turn be characterized by the treewidth (Definition 7) of that graph. Similar idea has also been exploited in markov2008simulating, in which the time complexity of tensor contraction (a special case of the Einsum operation) is characterized by the treewidth of the underlying tensor network.
@@
Definition 10 (Decomposition Graph).
@@
  • $V_{\mathcal{A}}=\textsc{Indices}(\mathcal{A})$,
@@
Example 2 (Einsum notations and its decomposition graphs).
@@
Figure 1: Examples of decomposition graphs of different decomposition signatures.
[…]
@@
Let $\mathcal{A}$ be an Einsum notation with no output, and $G_{\mathcal{A}}$ be the corresponding decomposition graph of $\mathcal{A}$. Then there exists an algorithm with a particular Einsum ordering such that for any set of tensors $\mathcal{T}$ that can be represented by the Einsum notation $\mathcal{A}$, the time complexity of $\textsc{Einsum}(\mathcal{T},\mathcal{A})$ is $O(|\mathcal{A}|n^{\mathsf{tw}(G_{\mathcal{A}})+1})$.
Remark 3 (Einsum summation ordering in practice).
We say that the time complexity given in the above proposition is only an optimistic estimate, because although it can be achieved with the optimal Einsum ordering corresponding to the treewidth (Definition 7), this particular ordering is usually difficult to find in practice. arnborg1987complexity have shown that finding the ordering that gives rise to the treewidth is an NP-complete problem. While it is ideal to evaluate $\textsc{Einsum}(\mathcal{T},\mathcal{A})$ in Algorithm 2 using this optimal ordering, we decide to directly plug in the Python library opt-einsum (daniel2018opt) to find the ordering in practice, because such a preexisting library has been optimized and stabilized over the years. opt-einsum offers both heuristic strategies such as greedy algorithms or exhaustive search such as dynamic programming, but the latter can be time-consuming when dealing with large-scale problems. As mentioned in the Introduction, dynamic programming is also used in zhang2025adaptive for a special class of $U$-statistics. Since the ordering only depends on the MD structure of the kernel $h$, for commonly encountered kernels, one could in principle find the optimal ordering via exhaustive search without accessing the data sample and re-use this ordering whenever the same type of kernel appears.
@@
3.3 The Algorithm of Computing $U$-Statistics